OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (2,13) and y-intercept (0,5). Write your answer in this form: (x1,y1),(x2,y2). If necessary, round to the nearest hundredth. @eragondragon666 @jlg030597 @agent0smith

OpenStudy (jlg030597):

is this algebra 2?

OpenStudy (anonymous):

No, Algebra 1.

OpenStudy (jlg030597):

wheres the picture just wondering I can help now that i know what type of math it is

OpenStudy (anonymous):

There is no picture =/..

OpenStudy (jlg030597):

hmmmmmm ok one second

OpenStudy (anonymous):

This is all there is..

OpenStudy (jlg030597):

whats the K stand for?

OpenStudy (jlg030597):

and the H?

OpenStudy (anonymous):

I don't know? Lol. New to this..

OpenStudy (jlg030597):

I mean ill help you.

OpenStudy (jlg030597):

Unless agent can help me and you understand.

OpenStudy (agent0smith):

The vertex is (h, k). Your vertex is (2, 13), plug those numbers in.

OpenStudy (jlg030597):

SO thatd be (y-13)=0(x-2)^2 right

OpenStudy (anonymous):

I'm guessing because that (y-k)=a(x-h)^2 You should be right? @agent0smith is he?

OpenStudy (agent0smith):

(y-13)=0(x-2)^2 Yes, but... Where did the 0 come from @jlg030597 ??

OpenStudy (jlg030597):

look at the picture she posted it confused me

OpenStudy (agent0smith):

Leave a as a.

OpenStudy (jlg030597):

(y-13)=a(x-2)^2

OpenStudy (agent0smith):

Now use the other point you're given, and plug in the x and y values for that point.

OpenStudy (jlg030597):

so (5-13)=a(0-2)^2

OpenStudy (jlg030597):

so then youd get -2=a-4

OpenStudy (agent0smith):

(-2)^2 is not -4

OpenStudy (agent0smith):

and 5-13 is not -2...

OpenStudy (jlg030597):

ops then -2=a4

OpenStudy (jlg030597):

right?

OpenStudy (agent0smith):

See my last post :P

OpenStudy (jlg030597):

Then what would it be?

OpenStudy (jlg030597):

-8=a4

OpenStudy (agent0smith):

...see my 2nd last post.

OpenStudy (jlg030597):

coming out as -4=a

OpenStudy (agent0smith):

start over from here (5-13)=a(0-2)^2

OpenStudy (jlg030597):

ok -8=a(-4)

OpenStudy (jlg030597):

5-13=-8 0-4=-4 so itd come out as 2=a

OpenStudy (agent0smith):

You made the same mistake i pointed out earlier

OpenStudy (jlg030597):

what ?

OpenStudy (agent0smith):

(-2)^2 is not -4

OpenStudy (jlg030597):

then itd still come out the opposite -2=a

OpenStudy (agent0smith):

Yep. Now you have a, h, and k, so you have the whole equation. Plug in y=0 to solve for the x intercepts.

OpenStudy (jlg030597):

(0-13)=-2(x-2)^2? sorry

OpenStudy (agent0smith):

Yes, solve for x.

OpenStudy (anonymous):

Won't it be a decimal? Considering 13 is an odd number?

OpenStudy (jlg030597):

-13=-2(x-2)^2 -13=-2(x2-4) is this right so far?

OpenStudy (agent0smith):

No, divide both sides by -2 first THEN take the square root of both sides btw you CANNOT distribute a power of 2 into brackets like you did there. Big mistake. (x-2)^2 = (x-2)(x-2) = x^2 - 4x + 4 NOT EVER EVER EVER (x^2 - 4)

OpenStudy (jlg030597):

Oh ops sorry im out of it right now can you help her im starting to get so confused and lost.

OpenStudy (anonymous):

So wait I think I got x.. Would it = 4.55? Evidence and my philosophy of it all.. (0-13)=-2(x-2)^2 so let's break it down .... 4.55-2 = 2.55 That gets rid of (x-2) .. So now we have -2(2.5495)^2 So what I did was 2.55^2 which = 6.50. That got rid of (2.55)^2 so now we got left: -2(6.50) which = -13. Is this correct??

OpenStudy (agent0smith):

You have an odd way of going about algebra... do things in the correct order, you appeared to do some kinda trial and error (0-13)=-2(x-2)^2 divide both sides by -2 first THEN take the square root of both sides. Don't forget to add a plus or minus sign.

OpenStudy (anonymous):

Okay... so.. (0-13)=-2(x-2)^2 ... Divide both sides FIRST by -2.. Would that leave 6.50 = x^2 ?

OpenStudy (anonymous):

Cause a negative divided by a negative (-13) and (-2) finalizes to a positive number, correct? And dividing both sides, thus canceling -2<-(x-2<-)^2 Leaving 6.50=x^2?

OpenStudy (agent0smith):

No... you can't magically make the -2 inside the brackets disappear.

OpenStudy (anonymous):

Lol awh.. Well then, so It'd be 6.50 = (x-2)^2?

OpenStudy (agent0smith):

Yes now take sqrt of both sides

OpenStudy (anonymous):

Okay, so.. 13=x-4 Then subtract 4 from both sides so it would be 13-4 = 9 = x So therefore, 9=x?

OpenStudy (agent0smith):

6.50 = (x-2)^2 take square root of both sides. What is the square root of x^2? If you know that, then what is the square root of (x-2)^2?

OpenStudy (agent0smith):

Also, the square root of 6.5 is not 13. Square root does not mean multiply by 2.

OpenStudy (anonymous):

Oh.. That's right.. Uhm.. I don't know how to do square root of (x-2)^2.. Can you please tell me? The 6.5 square root is 2.55

OpenStudy (agent0smith):

Tell me the square root of x^2.

OpenStudy (anonymous):

I do not know? Less it's 1...

OpenStudy (agent0smith):

Tell me the square root of 3^2

OpenStudy (anonymous):

3?

OpenStudy (agent0smith):

Now use that to tell me the square root of x^2.

OpenStudy (anonymous):

4? Lol.

OpenStudy (agent0smith):

Square root of 3^2 is 3. Square root of x^2 is...

OpenStudy (agent0smith):

Square root of 5^2 is 5. Notice the pattern.

OpenStudy (anonymous):

OH, so X

OpenStudy (agent0smith):

So the square root of (x-2)^2 is...

OpenStudy (anonymous):

x-2

OpenStudy (agent0smith):

Now finish. 6.50 = (x-2)^2

OpenStudy (anonymous):

What? LOL.. Thought we got rid of ^2?

OpenStudy (agent0smith):

Finish w/o mistakes from there.

OpenStudy (anonymous):

Okay..?? 6.50 = x - 2 Add 2 to both sides so 8.50 = x

OpenStudy (agent0smith):

lol i JUST said finish w/o mistakes from there, and you made a mistake :P

OpenStudy (anonymous):

-_- this is annyoing. Lol.. I'm confused.

OpenStudy (agent0smith):

We already went over how to solve it from here, now do it... we had to sidetrack because you had to learn square roots first: 6.50 = (x-2)^2

OpenStudy (anonymous):

That's what I did? You said "x^2" would = x... So this would be "6.50 = (x-2)"...?

OpenStudy (agent0smith):

Whatever you do to one side, you must do to the other side.

OpenStudy (anonymous):

Okay, so it would be 13 = x - 2 ?

OpenStudy (agent0smith):

You already forgot that taking a square root does not mean multiplying by 2...

OpenStudy (anonymous):

... Okay.. So.. 3.25 = x-2

OpenStudy (agent0smith):

Taking a square root also does not mean dividing by 2. You already found the square root earlier...

OpenStudy (anonymous):

.... Ok... 2.55...... so 2.55 = x - 2.....

OpenStudy (agent0smith):

And you also forgot that taking a square root leaves a plus or minus sign. \[\large \pm 2.55 = x - 2\]now solve for x you have to solve once for the + and once for the -), there will be two solutions.

OpenStudy (anonymous):

Ok... so... add two to both sides.. one for the plus..... so... -2.55 = x... -2.55 + 2... = -0.55 = x.... Then 2.55 + 2 = 4.55 = x... So -0.55, 4.55??

OpenStudy (agent0smith):

yes

OpenStudy (anonymous):

Thanks.. So that's the final answer? -0.55,4.55?

OpenStudy (agent0smith):

Have to write as coordinates, those are the x's, the y's are both zero

OpenStudy (anonymous):

So (-0.55,0),(4.55,0)..?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

Thanks.